OpenStudy (anonymous):

3. Which set of lengths could be the side lengths of a right triangle? (1 point) A) 20 cm, 22 cm, 29 cm B) 10 ft, 12 ft, 15 ft C) 7 km, 24 km, 28 km D) 13 in, 84 in, 85 in 4. A right triangle has a side length that measures 4 m and a hypotenuse that measures 8.5 m. What is the measure of the other side of the triangle? (1 point) A) 7.5 m B) 8.1 m C) 9.4 m D) 9.8 m

OpenStudy (anonymous):

5. You want to divide a square piece of paper into two equivalent triangles. If the square measures 20 cm on each side, what will the third side of each triangle measure? (1 point) A) 8.9 cm B) 20 cm C) 28.3 cm D) 40 cm

OpenStudy (anonymous):

For the first question, the easiest way to do this is to fill them all in to a^2 + b^2 = c^2. If a^2+b^2 don't add up to c^2, then it isn't your answer. Example: 10^2 + 12^2 = 15^2 100 + 144 = 225. 244 does not equal 225.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

The second question use the same equation. a^2 + b^2 = c^2. 4^2 + b^2 = 8.5^2 16 + b^2 = 72.25, subtract 16, b^2 = 56.25. b = \[\sqrt{56.25}\] which is also known as 7.5

OpenStudy (anonymous):

Final question, again your using a^2 + b^2 = c^2. The two lengths of each triangle = 20. 20^2+ 20^2 = c^2. 400 + 400 = c^2. 800 = c^2. C = \[\sqrt{800}\] or 28. 3

OpenStudy (anonymous):

so whats the first one?